OpenStudy (anonymous):

If no digit appears more than once, how many 3 digit numbers can be formed from the digits 2,4,5,7,9

OpenStudy (anonymous):

So how many different ways can you permute 3 numbers from a set of 5?

OpenStudy (anonymous):

does order matter? if not than its a combination

OpenStudy (anonymous):

Well it asks for 3 digit numbers, so I think the order must matter because 247 is a different 3 digit number than 274

OpenStudy (anonymous):

if 2,4,5 is the same as 2,5,4 then its \[\left(\begin{matrix}5 \\ 3\end{matrix}\right)\]

OpenStudy (anonymous):

all it said is if no digit appeared more than once, I'm assuming order doesn't matter since 2,5,4 is the same as 5,2,4 bc each digit appears only once in both cases

OpenStudy (amistre64):

without replacement is my take on it

OpenStudy (anonymous):

"how many 3 digit numbers can be formed"

OpenStudy (amistre64):

the logic goes, for me at least: there are 5 ways to pick the first digit there are 4 ways left to pick the second digit and there are 3 ways left to pick the last digit

OpenStudy (anonymous):

It's a permutation without repetition problem.

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

yeah it's a permutation

OpenStudy (anonymous):

Isn't it 5!/2!3!

OpenStudy (anonymous):

That's combinations. Permutations is 5!/(5-3)!

OpenStudy (anonymous):

no thats a combination

OpenStudy (anonymous):

5*4*3= 60

OpenStudy (anonymous):

I tend to mix them up, lol